# EIDSeg: A Pixel-Level Semantic Segmentation Dataset for Post-Earthquake Damage Assessment from Social Media Images

### v1

@@ Abstract
Code and data: https://github.com/HUILIHUANG413/EIDSeg
@@ Introduction
Earthquakes cause an estimated \$100 billion in damage globally each year, posing a persistent challenge to the resilience of infrastructure and the response to disasters (?). Rapid post-disaster analysis and response are the keys to rescue efforts, resource deployment, and reducing economic loss in the immediate aftermath of a major earthquake. However, damage assessment efforts often face a “cold start” problem: getting the analysis off the ground is delayed by challenges such as limited image availability, unpredictable weather conditions, and the need for labor intensive expert assessments, to name but a few.
@@
Conventional data collection is based on remote sensing imagery such as satellites and drones, combined with expert interpretation to delineate damaged areas and estimate severity (?; ?). Although widely adopted, these approaches face several well-documented limitations. First, acquiring high-resolution imagery is expensive and is often obstructed by weather conditions like clouds, fog, or rain, particularly in the critical early stages of a disaster (?; ?; ?). Second, damage analysis based on expert annotation is labor-intensive and requires substantial domain knowledge, making the process both time-consuming and resource-demanding. Recent studies have proposed automated tools for early disaster analysis, such as NASA’s Damage Proxy Maps (DPM) (?) and Microsoft’s Damage Model (?), but they typically return only a binary label that indicates whether an area is damaged, without describing the category of infrastructure or severity level.
@@
In contrast, mobile phones and social media platforms now provide an additional ground-level stream of photographs that are collected within minutes of an event. Social media platforms like X (formerly Twitter) expose street-scale damage invisible to aerial images. Despite this potential, social media imagery remains underutilized due to the lack of task-specific datasets with pixel-level annotations. While datasets such as EID (?), MEDIC (?), and DAD (?) offer multiclass labels for post-disaster damage, they are limited to image-level classification. Zhang et al. proposed the Damage Semantic Segmentation (DSS) dataset (?), but it only contains 607 post-earthquake social media images without benchmark results. As a result, there are currently no established standards for semantic segmentation of ground-level post-earthquake imagery (?).
@@
Developing a pixel-level segmentation benchmark for post-earthquake social media imagery presents two primary challenges. First, images vary widely in aspect ratio, resolution, and camera angle. Although annotators can typically assign a damage label to the most prominent building in the foreground, thoroughly segmenting all visible infrastructures, including distant or partially obstructed ones, requires significant manual effort. Second, neither existing standards in civil engineering for ground-level images nor semantic segmentation guidelines for aerial imagery are well suited to label infrastructure damage in ground-level social media images. Specifically, frameworks such as EMS-98 (?) and HAZUS (?) evaluate structural damage based on construction type (e.g., masonry, reinforced concrete), which is often unrecognizable in user-generated content. Prior aerial image segmentation datasets such as EarthquakeNet(drone-based) (?) and xBD(satellite-based) (?) offer semantic segmentation annotations; however, their aerial perspectives do not apply directly to ground-level perspectives (?). These limitations underscore the need for an annotation protocol and dataset specifically designed for social media imagery, which is both scalable and accessible to non-expert annotators working under high-pressure early-stage post-earthquake scenarios.
@@
To fill this gap, we introduce EIDSeg, the first large-scale semantic segmentation dataset of post-earthquake social media images. We curate the dataset by filtering low-quality images from the EID dataset (?) and integrating high-quality samples from the DSS dataset (?). To ensure our guideline is clear and accessible, six non-expert annotators with no civil engineering backgrounds labeled the dataset under the supervision of domain experts in post-earthquake engineering and infrastructure assessment. A three-phase quality control process was applied to enhance label consistency and reduce subjectivity.
@@
The inter-annotator analysis shows that the annotators achieved agreement over 70% in the final stage of the annotation, validating that the annotation guideline that we offer is clear and reproducible. We benchmark several state-of-the-art semantic segmentation models, including DeepLabV3+ (?), SegFormer (?), Mask2Former (?), BEiT (?), OneFormer (?), and EoMT (Encoder-only Mask Transformer) (?). Of these, the EoMT achieves the best performance, with an mIoU of 80.8% and a pixel accuracy of 90.3%.
@@
Civil engineering experts rely on drone images and field visits to post-earthquake sites to assess structural damage of infrastructures. A widely used standard, the European Macroseismic Scale (EMS-98) (?), defines five damage levels for buildings. G1 (Negligible to slight), G2 (Moderate), G3 (Substantial to heavy), G4 (Very heavy), and G5 (Destruction). The manual includes illustrative sketches and example images to guide human assessors. However, EMS-98 is limited in scope; it focuses solely on reinforced concrete and masonry buildings, excluding other types of infrastructure, such as transportation networks. FEMA’s HAZUS  (?) framework provides a more comprehensive classification system. It outlines 16 structural damage standards across 36 building categories, covering materials like wood, steel, and concrete. HAZUS also extends its damage evaluation to transportation and utility systems, using four severity levels: slight, moderate, extensive, and complete.
@@
xBD (?) Satellite 2019 $\times1024$ Instance Seg. / Cls. 192 4
TUE-CD (?) Satellite 2024 $\times256$ Semantic Seg. 2338 2
TE23D (?) Satellite 2025 $\times256$ Semantic Seg. 1,183 2
@@
MEDIC  (?) Social Media 2021 Varies Multi-Task Cls. 15,095 3
EID (?) Social Media 2025 Varies Cls. 13,513 4
DSS (?) Social Media 2025 Varies Semantic Seg. 607 3
EIDSeg(Ours) Social Media 2025 Varies Semantic Seg.n 3,266 5
Table 1: Overview of publicly available earthquake-damage datasets. ”Seg. = Segmentation”, ”Cls.” = classification, ”Post-EQ” = post-earthquake.
@@
Table 1 shows an overview of the existing data for the assessment of post-earthquake damage. Most existing datasets for earthquake damage segmentation are either part of broader multi-disaster collections or limited to a single geographic event. For example, the widely used xBD dataset (?) covers 19 disaster events, comprising 22,068 satellite images and 850,736 annotated building polygons. However, it includes only 192 pre-event and 192 post-event satellite images from the 2017 Mexico City earthquake, which is insufficient for large-scale earthquake-specific analysis. More recent segmentation datasets, such as TUE-CD (?) and TE23D (?), focus specifically on the 2023 Türkiye earthquakes but are geographically constrained, limiting their generalizability to other seismic regions.
Several datasets focus on the drone imagery. For instance, EarthquakeNet (?) provides high-resolution UAV-based images from Baoxing County with detailed annotations but includes only 69 images. Other datasets, such as AIDER (?) and AIDERv2 (?), include earthquake-related classes such as collapsed buildings to support the prediction of emergency response. However, these datasets lack damage severity labels, and the source events of the images are not specified, which limits their utility for systematic damage analysis.
For social media imagery, most studies have focused on image-level classification of damage severity (?; ?; ?; ?; ?; ?). Huang et al. developed the Earthquake Infrastructure Damage (EID) dataset, offering detailed labeling guidelines tailored for non-expert annotators and achieving strong performance in disaster analysis tasks (?). For segmentation, Zhang et al. introduced the DSS dataset, which provides pixel-level segmentation masks tailored for post-earthquake damage assessment (?). However, DSS lacks a formal benchmark and contains only 607 annotated images, limiting its generalization potential.
@@
Figure 1: Annotation workflow of EIDSeg. (a): Dataset creation: images are collected and filtered from the EID and DSS datasets, then independently labeled by annotation groups (A–C). Experts evaluate label quality, and high-quality samples are curated into the final EIDSeg dataset. (b): Label evaluation: Annotations undergo experts review following an ”Undesignated Checking” step, which address uncertain labels provided by annotators. If inter-annotator agreement (mIoU) exceeds a threshold $T$, the image proceeds to expert adjudication; otherwise, it is discarded to ensure dataset integrity.
@@
EIDSeg integrates images from both the EID dataset (?) and the DSS dataset (?). It comprises images from nine major earthquakes that span 2008 to 2023, enhancing the generalizability of the dataset in diverse earthquake scenarios.
The EID dataset compiles social media images from multiple sources, applying strict quality control and a well-documented annotation protocol. It covers five major events: the 2015 Nepal, 2015 Illapel, 2016 Ecuador, 2017 Mexico, and 2017 Iran–Iraq earthquakes. Each image is categorized into one of four classes: Irrelevant or Non-informative (DS0), No Damage (DS1), Mild Damage (DS2), and Severe Damage (DS3). Compared to earlier image-level datasets such as MEDIC (?), DAD (?), and CrisisBenchmark (?), EID provides finer class granularity and higher annotation quality. For EIDSeg, we include only DS1 to DS3 images, since DS0 images are mostly advertisements, maps, or logos that lack relevant infrastructure content.
The DSS dataset adds 607 images drawn from the 2008 Wenchuan, 2010 Haiti, 2015 Nepal, 2023 Morocco, and 2023 Türkiye earthquake, with additional examples retrieved through a Google search for ”earthquake damage”. DSS assigns three segmentation labels: Undamaged Structure, Damaged Structure, and Debris.
@@
Figure 3: Weekly segmentation agreement across annotators. Left: Inter-annotator mIoU (%) among Groups A, B, and C. Right: Average annotator–expert mIoU (%) within each group over time. Note that some annotators completed the task earlier due to graduation, which explains the missing data in later weeks (Group A in Week 8, and Group B in Weeks 7–8).
Referring to previous civil engineering manuals and remote sensing studies (?; ?; ?; ?), the EIDSeg dataset focuses on two common infrastructure types frequently observed in social media imagery: buildings and roads.
@@
  • Undamaged Road Structure (UD_Road): Roads that appear to be free of visible cracks, potholes, or major surface disruptions. Minor imperfections, such as hairline cracks (e.g., less than 1/8 inch (?)), are acceptable if they do not compromise structural integrity or user safety.
@@
Note that the EIDSeg damage scale defines only two categories for roads (undamaged and damaged) to align with what is visually structurally significant in post-disaster imagery. As shown in the rightmost column of Figure 2, roads generally do not undergo complete structural collapse as buildings do, which often results in clearly visible rubble or debris. Instead, road damage often appears as partial deterioration (?). Additionally, the presence of a ”debris-like” object on a roadway, such as collapsed structures or landslide material, does not necessarily imply that the road itself is damaged. In many cases, such rubbles originate from surrounding terrain (e.g., hillsides or mountains), rather than from the road infrastructure itself, and are therefore not included in our structural damage categories.
In addition to the structural damage classes, we introduce an Undesignated class to reduce annotation ambiguity and ease the labeling burden in challenging cases. To maintain consistency and quality, all Undesignated instances are reviewed by domain experts during each annotation round, allowing for a careful and conservative judgment. This class is excluded from the final EIDSeg dataset statistics. For all other annotation procedures, we follow the VOC2011 segmentation guidelines (?).
@@
To ensure annotation quality, the experts prepared a curated set of 100 gold standard images. For each annotation round, they randomly selected 20–25 gold images and embedded them in the labeling batches assigned to each group. Two agreement metrics were tracked: (i) annotator-vs-expert mIoU (on gold images), and (ii) inter-annotator mIoU (on jointly labeled images). Annotators with expert agreement below 0.60 were flagged for retraining. Images with inter-annotator agreement less than 0.60 were excluded from the final dataset to avoid the subjectivity issue  (?). The $0.60$ threshold was selected based on the highest annotator-expert agreement observed during Phase 1 training. This ensured that only consistent annotations were carried forward to Phase 3.
@@
Figure 2 presents several representative examples from the EIDSeg dataset, along with their corresponding color-coded annotation masks. Figure 3 illustrates the average annotator-vs-expert mIoU and inter-annotator mIoU for each group over time during Phase 2. The agreement remained consistently high throughout the annotation process and exhibited an upward trend, indicating improved annotator consistency and better alignment with the expert-labeled ground truth over time. By the final round, all groups achieved agreement above 70%, a promising outcome for disaster assessment tasks. For comparison, the image-level EID dataset reached 86.6% agreement in its final round (?), while EIDSeg demands more complex pixel-level labeling across finer-grained classes.
Figure 4 displays the pixel distribution in EIDSeg. In particular, the dataset contains a higher proportion of pixels labeled damaged compared to undamaged. This trend reflects the underlying class distribution of the EID image corpus, which contains more disaster-affected scenes (52.1%) than undamaged ones (26.86%) (?). This suggests that social media users are more likely to capture and share images of visible damage than undisturbed scenes after an earthquake.
Figure 4: Pixel distribution across segmentation classes in the EIDSeg dataset. Each bar shows the total pixel count per class, along with the percentage it contributes to the full dataset. Class colors are chosen to reflect semantic segmentation label: Undamaged Building to Damaged Building to Debris; and Undamaged Road to Damaged Road.
@@
DeepLabV3 (?) ResNet‑101 Cityscapes $512^{2}$ $67.1$ $68.2$ $80.6$ $79.29$ $58.75$
SegFormer (?) MiT‑B5 Cityscapes $512^{2}$ $74.4$ $75.2$ $86.9$ $110.16$ $84.60$
Mask2Former (?) Swin‑S Cityscapes $512^{2}$ $76.1$ $77.1$ $87.5$ $93.21$ $68.72$
@@
OneFormer (?) Swin‑L Cityscapes $512^{2}$ $79.2$ $78.3$ $88.2$ $1042.14$ $218.77$
EoMT (?) ViT‑L Cityscapes $1024^{2}$ $80.8$ $80.9$ $90.3$ $1341.85$ $319.02$
@@
Dataset Preparation: We follow the data splitting strategy adopted in the xBD paper (?) and divide the EIDSeg dataset into 80% for training, 10% for validation and 10% for testing.
Models and Training: We evaluated eight state-of-the-art semantic segmentation models: DeepLabV3+(ResNet-101) (?), SegFormer (B5) (?), Mask2Former (Swin-Base and Swin-L) (?), BEiT (ViT-B and ViT-L) (?), OneFormer (Swin-L) (?), and EoMT (ViT-L) (?). All models are initialized with Cityscapes-pre-trained weights except for BEiT, which uses the ADE20K-pre-trained checkpoint due to the absence of a Cityscapes version.
@@
As discussed in Related Work, the DSS dataset is the only post-earthquake social media dataset for semantic segmentation. The original DSS dataset reported only one baseline: SegFormer‑B5 trained for three-class segmentation, which achieved an mIoU of 72% (?). In contrast, Table 2 presents a comprehensive evaluation of eight modern architectures on the proposed EIDSeg benchmark, which adopts a more fine-grained five-class labeling. Under these conditions, SegFormer‑B5 attains an mIoU of 74.4%, yielding a 2.4% improvement over the DSS result while addressing a more detailed five-class labeling that includes two additional categories. This gain highlights the advantages of our expanded image corpus and refined annotation strategy, which collectively improve class balance and reduce labeling noise.
@@
Table 3 presents the class-wise IoU for each model. Among all classes, Undamaged Building and Damaged Building consistently exhibit lower IoU scores. This can be attributed to two main factors. First, during the annotation process for EIDSeg, the annotators followed a conservative guideline: if a building appeared structurally intact but exhibited visible cracks not attributable to wear and tear, it was labeled as Damaged Building due to lack of information about its interior condition. Although this labeling strategy supports cautious decision-making in post-disaster analysis, it may lead to confusion for models when encountering visually similar undamaged structures. Second, since the dataset is collected from social media, buildings in the background often appear blurry or lack sufficient detail, making it difficult to assess structural damage. This reflects an inherent trade-off in the use of social media imagery. Although it enables access to timely data from disaster-stricken areas where professionals may not be present, it also limits control over image quality and resolution. Visual examples of misclassified images are provided in Appendix B.

### v2

@@ Abstract
Code and Datasetshttps://github.com/HUILIHUANG413/EIDSeg
Extended versionhttps://arxiv.org/abs/2511.06456
@@ Introduction
Earthquakes cause an estimated \$100 billion in damage globally each year, posing a persistent challenge to the resilience of infrastructure and the response to disasters (GEM2023seismic). Rapid post-disaster analysis and response are the keys to rescue efforts, resource deployment, and reducing economic loss in the immediate aftermath of a major earthquake. However, damage assessment efforts often face a “cold start” problem: getting the analysis off the ground is delayed by challenges such as limited image availability, unpredictable weather conditions, and the need for labor intensive expert assessments, amongst others.
@@
Conventional data collection is based on remote sensing imagery such as satellites and drones, combined with expert interpretation to delineate damaged areas and estimate severity (Gupta2019xBD; Huang2025Fidelity). Although widely adopted, these approaches face several well-documented limitations. First, acquiring high-resolution imagery is expensive and is often obstructed by weather conditions like clouds, fog, or rain, particularly in the critical early stages of a disaster (Gupta2019xBD; freddi2021innovations; drones8110638). Second, damage analysis based on expert annotation is labor-intensive and requires substantial domain knowledge, making the process both time-consuming and resource-demanding. Recent studies have proposed automated tools for early disaster analysis, such as NASA’s Damage Proxy Maps (DPM) (NASA2020ARIA_DPM) and Microsoft’s Damage Model (Microsoft2023TurkeyEarthquake), but they typically return only a binary label that indicates whether an area is damaged, without describing the category of infrastructure or severity level.
@@
In contrast, mobile phones and social media platforms now provide an additional ground-level stream of photographs that are collected within minutes of an event. Social media platforms like X (formerly Twitter) expose street-scale damage invisible to aerial images. Despite this potential, social media imagery remains underutilized due to the lack of task-specific datasets with pixel-level annotations. While datasets such as EID (Huang2025Fidelity), MEDIC (Alam2021MEDIC), and DAD (DAD-Dataset) offer multiclass labels for post-disaster damage, they are limited to image-level classification. Zhang et al. proposed the Damage Semantic Segmentation (DSS) dataset (Zhang2025PixelsDamageSeverity), but it only contains 607 post-earthquake social media images without benchmark results. As a result, there are currently no established standards for semantic segmentation of ground-level post-earthquake imagery (Zhu2020MSNet).
@@
Developing a pixel-level segmentation benchmark for post-earthquake social media imagery presents two primary challenges. First, images vary widely in aspect ratio, resolution, and camera angle. Although annotators can typically assign a damage label to the most prominent building in the foreground, thoroughly segmenting all visible infrastructures, including distant or partially obstructed ones, requires significant manual effort. Second, neither existing standards in civil engineering for ground-level images nor semantic segmentation guidelines for aerial imagery are well suited to label infrastructure damage in ground-level social media images. Specifically, frameworks such as EMS-98 (Grunthal1998EMS98) and HAZUS (FEMA2020HAZUS4_2_SP3) evaluate structural damage based on construction type (e.g., masonry, reinforced concrete), which is often unrecognizable in user-generated content. Prior aerial image segmentation datasets such as EarthquakeNet(drone-based) (Jiang2024EarthquakeNet) and xBD(satellite-based) (Gupta2019xBD) offer semantic segmentation annotations; however, their aerial perspectives do not apply directly to ground-level perspectives (Huang2025Fidelity). These limitations underscore the need for an annotation protocol and dataset specifically designed for social media imagery, which is both scalable and accessible to non-expert annotators working under high-pressure early-stage post-earthquake scenarios.
@@
To fill this gap, we introduce EIDSeg, the first large-scale semantic segmentation dataset of post-earthquake social media images. We curate the dataset by filtering low-quality images from the EID dataset (Huang2025Fidelity) and integrating high-quality samples from the DSS dataset (Zhang2025PixelsDamageSeverity). To ensure our guideline is clear and accessible, six non-expert annotators with no civil engineering backgrounds labeled the dataset under the supervision of domain experts in post-earthquake engineering and infrastructure assessment. A three-phase quality control process was applied to enhance label consistency and reduce subjectivity.
@@
The inter-annotator analysis shows that the annotators achieved agreement over 70% in the final stage of the annotation, validating that the annotation guideline that we offer is clear and reproducible. We benchmark several state-of-the-art semantic segmentation models, including DeepLabV3+ (Chen2018DeepLabV3Plus), SegFormer (Xie2021SegFormer), Mask2Former (Cheng2022Mask2Former), BEiT (Bao2022BEiT), OneFormer (Jain2023OneFormer), and EoMT (Encoder-only Mask Transformer) (Kerssies2025EoMT). Of these, the EoMT achieves the best performance, with an mIoU of 80.8% and a pixel accuracy of 90.3%.
@@
Civil engineering experts rely on drone images and field visits to post-earthquake sites to assess structural damage of infrastructures. A widely used standard, the European Macroseismic Scale (EMS-98) (Grunthal1998EMS98), defines five damage levels for buildings. G1 (Negligible to slight), G2 (Moderate), G3 (Substantial to heavy), G4 (Very heavy), and G5 (Destruction). The manual includes illustrative sketches and example images to guide human assessors. However, EMS-98 is limited in scope; it focuses solely on reinforced concrete and masonry buildings, excluding other types of infrastructure, such as transportation networks. FEMA’s HAZUS  (FEMA2020HAZUS4_2_SP3) framework provides a more comprehensive classification system. It outlines 16 structural damage standards across 36 building categories, covering materials like wood, steel, and concrete. HAZUS also extends its damage evaluation to transportation and utility systems, using four severity levels: slight, moderate, extensive, and complete.
@@
xBD (Gupta2019xBD) Satellite 2019 $\times1024$ Instance Seg. / Cls. 192 4
TUE-CD (Liu2024BuildingChangeDetectionTUECD) Satellite 2024 $\times256$ Semantic Seg. 2338 2
TE23D (Ekkazan2025TE23D) Satellite 2025 $\times256$ Semantic Seg. 1,183 2
@@
MEDIC  (Alam2021MEDIC) Social Media 2021 Varies Multi-Task Cls. 15,095 3
EID (Huang2025Fidelity) Social Media 2025 Varies Cls. 13,513 4
DSS (Zhang2025PixelsDamageSeverity) Social Media 2025 Varies Semantic Seg. 607 3
EIDSeg(Ours) Social Media 2025 Varies Semantic Seg. 3,266 5
Table 1: Overview of publicly available earthquake-damage datasets. “Seg. = Segmentation”, “Cls.” = classification, “Post-EQ” = post-earthquake.
@@
Table 1 shows an overview of the existing data for the assessment of post-earthquake damage. Most existing datasets for earthquake damage segmentation are either part of broader multi-disaster collections or limited to a single geographic event. For example, the widely used xBD dataset (Gupta2019xBD) covers 19 disaster events, comprising 22,068 satellite images and 850,736 annotated building polygons. However, it includes only 192 pre-event and 192 post-event satellite images from the 2017 Mexico City earthquake, which is insufficient for large-scale earthquake-specific analysis. More recent segmentation datasets, such as TUE-CD (Liu2024BuildingChangeDetectionTUECD) and TE23D (Ekkazan2025TE23D), focus specifically on the 2023 Türkiye earthquakes but are geographically constrained, limiting their generalizability to other seismic regions.
Several datasets focus on the drone imagery. For instance, EarthquakeNet (Jiang2024EarthquakeNet) provides high-resolution UAV-based images from Baoxing County with detailed annotations but includes only 69 images. Other datasets, such as AIDER (Kyrkou2019AIDER) and AIDERv2 (Shianios2023AIDERv2), include earthquake-related classes such as collapsed buildings to support the prediction of emergency response. However, these datasets lack damage severity labels, and the source events of the images are not specified, which limits their utility for systematic damage analysis.
For social media imagery, most studies have focused on image-level classification of damage severity (DAD-Dataset; Alam2017Image4ActOS; Nguyen2017AutomaticIF; Nia2017BuildingDA; Alam2021MEDIC; crisisdataset2020icwsm). Huang et al. developed the Earthquake Infrastructure Damage (EID) dataset, offering detailed labeling guidelines tailored for non-expert annotators and achieving strong performance in disaster analysis tasks (Huang2025Fidelity). For segmentation, Zhang et al. introduced the DSS dataset, which provides pixel-level segmentation masks tailored for post-earthquake damage assessment (Zhang2025PixelsDamageSeverity). However, DSS lacks a formal benchmark and contains only 607 annotated images, limiting its generalization potential.
@@
Figure 1: Annotation workflow of EIDSeg. (a): Dataset creation: images are collected and filtered from the EID and DSS datasets, then independently labeled by annotation groups (A–C). Experts evaluate label quality, and high-quality samples are curated into the final EIDSeg dataset. (b): Label evaluation: Annotations undergo experts review following an “Undesignated Checking” step, which address uncertain labels provided by annotators. If inter-annotator agreement (mIoU) exceeds a threshold $T$, the image proceeds to expert adjudication; otherwise, it is discarded to ensure dataset integrity.
@@
EIDSeg integrates images from both the EID dataset (Huang2025Fidelity) and the DSS dataset (Zhang2025PixelsDamageSeverity). It comprises images from nine major earthquakes that span 2008 to 2023, enhancing the generalizability of the dataset in diverse earthquake scenarios.
The EID dataset compiles social media images from multiple sources, applying strict quality control and a well-documented annotation protocol. It covers five major events: the 2015 Nepal, 2015 Illapel, 2016 Ecuador, 2017 Mexico, and 2017 Iran–Iraq earthquakes. Each image is categorized into one of four classes: Irrelevant or Non-informative (DS0), No Damage (DS1), Mild Damage (DS2), and Severe Damage (DS3). Compared to earlier image-level datasets such as MEDIC (Alam2021MEDIC), DAD (DAD-Dataset), and CrisisBenchmark (crisisdataset2020icwsm), EID provides finer class granularity and higher annotation quality. For EIDSeg, we include only DS1 to DS3 images, since DS0 images are mostly advertisements, maps, or logos that lack relevant infrastructure content.
The DSS dataset adds 607 images drawn from the 2008 Wenchuan, 2010 Haiti, 2015 Nepal, 2023 Morocco, and 2023 Türkiye earthquake, with additional examples retrieved through a Google search for “earthquake damage”. DSS assigns three segmentation labels: Undamaged Structure, Damaged Structure, and Debris.
@@
Figure 3: Weekly segmentation agreement across annotators. Left: Inter-annotator mIoU (%) among Groups A, B, and C. Right: Average annotator–expert mIoU (%) within each group over time. Note that some annotators completed the assigned tasks earlier, which explains the different end dates in the figure (Group A in Week 8, and Group B in Weeks 7–8).
Referring to previous civil engineering manuals and remote sensing studies (Grunthal1998EMS98; FEMA2020HAZUS4_2_SP3; Jiang2024EarthquakeNet; Zhang2025PixelsDamageSeverity), the EIDSeg dataset focuses on two common infrastructure types frequently observed in social media imagery: buildings and roads.
@@
  • Undamaged Road Structure (UD_Road): Roads that appear to be free of visible cracks, potholes, or major surface disruptions. Minor imperfections, such as hairline cracks (e.g., less than 1/8 inch (FEMA2020HAZUS4_2_SP3)), are acceptable if they do not compromise structural integrity or user safety.
@@
Note that the EIDSeg damage scale defines only two categories for roads (undamaged and damaged) to align with what is visually structurally significant in post-disaster imagery. As shown in the rightmost column of Figure 2, roads generally do not undergo complete structural collapse as buildings do, which often results in clearly visible rubble or debris. Instead, road damage often appears as partial deterioration (Jiang2024EarthquakeNet). Additionally, the presence of a “debris-like” object on a roadway, such as collapsed structures or landslide material, does not necessarily imply that the road itself is damaged. In many cases, such rubbles originate from surrounding terrain (e.g., hillsides or mountains), rather than from the road infrastructure itself, and are therefore not included in our structural damage categories.
In addition to the structural damage classes, we introduce an Undesignated class to reduce annotation ambiguity and ease the labeling burden in challenging cases. To maintain consistency and quality, all Undesignated instances are reviewed by domain experts during each annotation round, allowing for a careful and conservative judgment. This class is excluded from the final EIDSeg dataset statistics. For all other annotation procedures, we follow the VOC2011 segmentation guidelines (Everingham2010VOC).
@@
To ensure annotation quality, the experts prepared a curated set of 100 “gold standard” images. For each annotation round, they randomly selected 20–25 gold images and embedded them in the labeling batches assigned to each group. Two agreement metrics were tracked: (i) annotator-vs-expert mIoU (on gold standard images), and (ii) inter-annotator mIoU (on jointly labeled images). Annotators with expert agreement below 0.60 were flagged for retraining. Images with inter-annotator agreement less than 0.60 were excluded from the final dataset to avoid the subjectivity issue  (Huang2025Fidelity). The $0.60$ threshold was selected based on the highest annotator-expert agreement observed during Phase 1 training. This ensured that only consistent annotations were carried forward to Phase 3.
@@
Figure 2 presents several representative examples from the EIDSeg dataset, along with their corresponding color-coded annotation masks. Figure 3 illustrates the average annotator-vs-expert mIoU and inter-annotator mIoU for each group over time during Phase 2. The agreement remained consistently high throughout the annotation process and exhibited an upward trend, indicating improved annotator consistency and better alignment with the expert-labeled ground truth over time. By the final round, all groups achieved agreement above 70%, a promising outcome for disaster assessment tasks. For comparison, the image-level EID dataset reached 86.6% agreement in its final round (Huang2025Fidelity), while EIDSeg demands more complex pixel-level labeling across finer-grained classes.
Figure 4 displays the pixel distribution in EIDSeg. In particular, the dataset contains a higher proportion of pixels labeled damaged compared to undamaged. This trend reflects the underlying class distribution of the EID image corpus, which contains more disaster-affected scenes (52.1%) than undamaged ones (26.86%) (Huang2025Fidelity). This suggests that social media users are more likely to capture and share images of visible damage than undisturbed scenes after an earthquake.
Figure 4: Pixel distribution across segmentation classes in the EIDSeg dataset. Each bar shows the total pixel count per class, along with the percentage it contributes to the full dataset. Class colors are chosen to reflect semantic segmentation label: Undamaged Building, Damaged Building, Debris; and Undamaged Road, Damaged Road.
@@
DeepLabV3 (Chen2018DeepLabV3Plus) ResNet‑101 Cityscapes $512^{2}$ $67.1$ $68.2$ $80.6$ $79.29$ $58.75$
SegFormer (Xie2021SegFormer) MiT‑B5 Cityscapes $512^{2}$ $74.4$ $75.2$ $86.9$ $110.16$ $84.60$
Mask2Former (Cheng2022Mask2Former) Swin‑S Cityscapes $512^{2}$ $76.1$ $77.1$ $87.5$ $93.21$ $68.72$
@@
OneFormer (Jain2023OneFormer) Swin‑L Cityscapes $512^{2}$ $79.2$ $78.3$ $88.2$ $1042.14$ $218.77$
EoMT (Kerssies2025EoMT) ViT‑L Cityscapes $1024^{2}$ $80.8$ $80.9$ $90.3$ $1341.85$ $319.02$
@@
Dataset Preparation: We follow the data splitting strategy adopted in the xBD paper (Gupta2019xBD) and divide the EIDSeg dataset into 80% for training, 10% for validation and 10% for testing.
Models and Training: We evaluated eight state-of-the-art semantic segmentation models: DeepLabV3+(ResNet-101) (Chen2018DeepLabV3Plus), SegFormer (B5) (Xie2021SegFormer), Mask2Former (Swin-Base and Swin-L) (Cheng2022Mask2Former), BEiT (ViT-B and ViT-L) (Bao2022BEiT), OneFormer (Swin-L) (Jain2023OneFormer), and EoMT (ViT-L) (Kerssies2025EoMT). All models are initialized with Cityscapes-pre-trained weights except for BEiT, which uses the ADE20K-pre-trained checkpoint due to the absence of a Cityscapes version.
@@
As discussed in Related Work, the DSS dataset is the only post-earthquake social media dataset for semantic segmentation. The original DSS dataset reported only one baseline: SegFormer‑B5 trained for three-class segmentation, which achieved an mIoU of 72% (Zhang2025PixelsDamageSeverity). In contrast, Table 2 presents a comprehensive evaluation of eight modern architectures on the proposed EIDSeg benchmark, which adopts a more fine-grained five-class labeling. Under these conditions, SegFormer‑B5 attains an mIoU of 74.4%, yielding a 2.4% improvement over the DSS result while addressing a more detailed five-class labeling that includes two additional categories. This gain highlights the advantages of our expanded image corpus and refined annotation strategy, which collectively improve class balance and reduce labeling noise.
@@
Table 3 presents the class-wise IoU for each model. Among all classes, Undamaged Building and Damaged Building consistently exhibit lower IoU scores. This can be attributed to two main factors. First, during the annotation process for EIDSeg, the annotators followed a conservative guideline: if a building appeared structurally intact but exhibited visible cracks not attributable to wear and tear, it was labeled as Damaged Building due to lack of information about its interior condition. Although this labeling strategy supports cautious decision-making in post-disaster analysis, it may lead to confusion for models when encountering visually similar undamaged structures. Second, since the dataset is collected from social media, buildings in the background often appear blurry or lack sufficient detail, making it difficult to assess structural damage. This reflects an inherent trade-off in the use of social media imagery. Although it enables access to timely data from disaster-stricken areas where professionals may not be present, it also limits control over image quality and resolution. Visual examples of misclassified images and the leave-one-event-out results are presented in Appendix B and Appendix D, respectively.